how to you get what number is between 3/4 and 2/3?
Answer by oberobic(2304)

If you wanted to find the number that is half the distance between 3/4 and 2/3, you could multiply the difference by 1/2 and add to the smaller fraction.
.
3/4 = 9/12
2/3 = 8/12
----------------subtract
3/4 -2/3 = 1/12
.
1/2 * 1/12 = 1/24
.
2/3 + 1/24 = 16/24 + 1/24 = 17/24
.
Done.